Previous incidents
API Maintenance
Resolved Mar 07 at 02:00pm GMT
During this timeframe the API might be rendered unresponsive for a brief period of time.
Emergency maintenance to update critical switching firmware
Resolved Mar 04 at 12:49am GMT
Service has been restored, we will provide affected customers with a more accurate rundown of the outage details as well as applicable compensation in the coming days.
4 previous updates
Metal deploy migration
Resolved Feb 21 at 01:30pm GMT
We are currently migrating some internal services to our new operations boxes in order to increase redundancy and reliability.
EDIT:
This migration was successfully completed at 13:53 GMT.
Emergency routing changes
Resolved Feb 17 at 03:30pm GMT
We will be performing a BGP session swap in London.
This should not be disruptive to clients.
Edit
There was a brief disruption to services in our London location at 23:20pm GMT this was due to a routing error with our upstream when we swapped our BGP session. This has been resolved and no further disruption is expected.